# Towards Personalized Treatment Plan: Geometrical Model-Agnostic Approach to Counterfactual Explanations

### v1

@@ Abstract
What if a person never smoked? Would they have lung cancer? What identifies those who don’t have cancer from those who don’t? Counterfactuals are statements of “alternative hypothesis” that describe the outcomes of an alternative scenario (Pearl & Mackenzie, 2018). They question the possibilities of a scenario that never occurred. While counterfactuals are hard to analyze, we can analyze counterfactual explanations which are more rigorously based. Counterfactual explanations define the “smallest” change to a data point such that the change flips the predicted outcome for a data point (Molnar, 2025). In our article, we describe a method for generating counterfactual explanations in high-dimensional spaces using four steps that involve fitting our dataset to a model, finding the decision boundary, determining constraints on the problem, and computing the closest point (counterfactual explanation) from that boundary. We propose a discretized approach where we find many discrete points on the boundary and then identify the closest feasible counterfactual explanation. This method, which we later call Segmented Sampling for Boundary Approximation (SSBA), applies binary search to find decision boundary points and then searches for the closest boundary point. Across four datasets of varying dimensionality, we show that our method can outperform current methods for counterfactual generation with reductions in distance between $5\%$ to $50\%$ in terms of the $L_{2}$ norm. Our method can also handle real-world constraints by restricting changes to immutable and categorical features, such as age, gender, sex, height, and other related characteristics such as the case for a health-based dataset. In terms of runtime, the SSBA algorithm generates decision boundary points on multiple orders of magnitude in the same given time when we compare to a grid-based approach. In general, our method provides a simple and effective model-agnostic method that can compute nearest feasible (i.e. realistic with constraints) counterfactual explanations. All of our results and our code can be found here at this link: https://github.com/ dsin85691/SSBA_For_Counterfactuals
@@ 1 Introduction
Imagine two scenarios: “If Joe had not smoked for $30$ years, he would have lived,” or “if Joe were 30 pounds lighter, he would not have heart disease.” While we may know that Joe smoked or was 30 pounds heavier, we may ask a question where we imagine outcomes to scenarios where the initial conditions were altered. These scenarios are among many described by Dr. Judea Pearl as “counterfactuals.” Counterfactuals are statements of “alternative hypothesis” that question the outcome given an alternative set of initial conditions (Pearl & Mackenzie, 2018). When compared to other types of causal relationships commonly studied in machine learning, such as association or intervention, counterfactuals are “problematic” for study since we deal with outcomes that are not easily identifiable from the dataset. When given observational data for a person, we can never observe more than one potential outcome for a given person (Pearl & Mackenzie, 2018). This makes it challenging to create logically consistent counterfactuals in a setting where we cannot perfectly replicate the same inputs. Standard machine learning methods, such as regression, can be helpful for questions of association where we observe how, for a set of variables $X$ and $Y$, a change in $X$ can alter the value of $Y$. With intervention, it is common to set up conditions so that other variables do not affect the primary variable of interest and its corresponding dependent variable. Randomized Control Trials (RCTs) were considered the “gold standard of causal analysis,” and it was common to randomize subjects into different groups to minimize the effect of confounding variables on treatment (Pearl & Mackenzie, 2018). However, with counterfactuals, we must estimate the relationships between variables in a world where “observed facts are bluntly negated” (Pearl & Mackenzie, 2018).
@@
To compute counterfactuals, Dr. Pearl relied on a method that made the use of structural causal models (SCMs) which are graphical models consisting of exogenous variables ($U$) and endogenous variables ($V$) and a “set of functions $f$ that assigns each variable in $V$ a value based on the other variables in the model” (Pearl et al., 2016). Dr. Pearl provides a three-step process for computing any deterministic counterfactual with a structural causal model $M$, which includes abduction, action, and prediction (Pearl et al., 2016). Abduction uses evidence $E=e$ to set the values of the exogenous variables in $M$. Action modifies the original model $M$ by removing the structural equations for the variables in $X$ and replacing them with the appropriate functions $X=x$ to obtain the model $M_{x}$ (Pearl et al., 2016). Prediction uses the modified $M_{x}$ and the value of exogenous variables in $U$ to compute the value of $Y$, the consequence of the counterfactual (Pearl et al., 2016). Similar to the idea used for Bayesian networks, we can use evidence or conditional information on the exogenous variables to simplify and construct a modified model that we can use to predict the outcome.
@@
In high-dimensional datasets, determining the structural causal model beforehand can be challenging. The use of causal discovery algorithms may require conditions such as causal sufficiency, which necessitates knowing all confounding variables within the dataset. Due to the difficulty of building a structural causal model and inferring causal relationships directly from the dataset, we address these issues by simplifying the problem to binary classification, assuming sufficient data for both classes, and relying on a formal definition of counterfactual explanations. As described by Dandl and Molnar [2025], a counterfactual explanation (CFE) describes “the smallest change to the feature values that changes the prediction to a predefined output” (Molnar, 2025). Given the direct causal relationship between input and prediction of a machine learning model, we can observe “scenarios in which the prediction changes in a relevant way, like a flip in predicted class” (Molnar, 2025). For CFEs, we find the smallest change across features such that the model flips a prediction value for a specific data point. For binary classification, such a flip would mean becoming a member of the other class as predicted by the model. In terms of criteria, we would like these explanations to be “as similar as possible to the instance regarding feature values” and require us to change as “few features as possible” (Molnar, 2025). Finally, we would want feature values that are likely for the counterfactuals (Molnar, 2025). Producing a CFE that generates illogical features, such as a younger person, or a person with an age greater than $150$, would not be realistic.
@@
To generate a CFE, we can employ multiple approaches, including randomly selecting counterfactuals, optimizing the choice of counterfactuals, and generating decision boundary points (our approach). One approach involves “randomly changing feature values of the instance of interest and stopping when the desired output is predicted” (Molnar, 2025). As another approach, Wachter et. al. [2017] proposed an objective loss function where the two terms of the objective reduce the quadratic distance between the model prediction for the counterfactual $\mathbf{x^{\prime}}$ and a second term to minimize the distance between the instance $\mathbf{x}$ and the counterfactual $\mathbf{x^{\prime}}$ (Molnar, 2025; Wachter et al., 2017). Across multiple iterations, we would find a counterfactual $x^{\prime}$ that minimizes the objective loss function, and then we return that counterfactual at the end.
@@
Dandl et. al. [2020] also implemented a four-objective loss function that minimized all of the above criteria for counterfactual explanations (Molnar, 2025). Similar to Wachter et. al. [2017], they applied an combined objective loss function to optimize an appropriate counterfactual $\mathbf{x^{\prime}}$ selectively. $\mathbf{o_{1}},\cdots\mathbf{o_{4}}$ are objective functions that each individually represents one of the criteria above for CFEs.
@@
Among model-agnostic methods developed, Mothilal, Sharma, and Tan implemented DiCE (Diverse Counterfactual Explanations for ML), a Python package that generates CFEs for both gradient-based and model-agnostic methods (Mothilal et al., 2020). Their gradient-based method optimizes over a combined loss function over all generated counterfactuals that take into account the difference between the machine predictions $f(c_{i})$ and the desired outcomes $y$ (first term), the distance between the counterfactual $c_{i}$ and the instance $x$, and the diversity among different counterfactuals.
@@
(Mothilal et al., 2020)
In their model-agnostic approaches, they adopt three different algorithms to generate a multitude of CFEs, which are their KDTrees algorithm, their random algorithm, and their genetic algorithm. The KDTree algorithm queries the $k$ closest counterfactuals from the dataset using KDTrees (Mothilal et al., 2020). The random algorithm generates a random set of counterfactuals that are generally in the opposite class relative to the original instance (Mothilal et al., 2020). Finally, the genetic algorithm tries to find the best counterfactuals close to the query point using a convergent algorithm that computes the best counterfactuals for a given generation based on a loss function taking into account criteria such as proximity and sparsity. It then generates a new generation based on $50\%$ of the fittest members of the current generation (Mothilal et al., 2020). All of these methods use the original instance $\mathbf{x}$ and the counterfactual $\mathbf{x^{\prime}}$ to iterate and find a new counterfactual that matches the criteria and reduce the objective loss function between the counterfactual and the original instance.
@@ 2 Related Works
In this article, we primarily examine our methods with respect to DiCE and Alibi, which are both open-source Python packages for high-dimensional counterfactual generation. Wexler et al. [2019] present a similar methodology for nearest counterfactual generation. However, their approach selects a data point of a differing target class within the dataset (Wexler et al., 2019). While their tool can generate nearest counterfactual explanations based on the dataset, their “What-If” tool is specialized for two dimensions and does not readily apply to higher dimensions. Karimi et. al. [2019] implemented MACE (model-agnostic counterfactual explanation) with the use of SMT solvers, but their work applies a distance metric based on $L_{0}$, $L_{1}$, and $L_{\infty}$ and not $L_{2}$ for $n\geq 2$ (Karimi et al., 2020). Additionally, the discontinuation of the MACE project has made it difficult to replicate results applying $L_{2}$ norms. Due to limited research on counterfactual generation using $L_{2}$, we limit our work with respect to DiCE and Alibi.
@@ 3 Methodology
Our methodology is based on a discretized approach to find the nearest counterfactual explanation, which focuses on generating decision boundary points rather than optimizing a counterfactual with an objective loss function.
@@
In the results section, we compare the computed counterfactual explanations and the runtime for various counterfactual generation methods which include DiCE, Alibi, and our SSBA method which we describe later. As discussed earlier, Mothital et al. released DiCE, an open-source repository that adopts much of the recent research in the space of generating CFEs. Additionally, Janis Klaise et. al. [2021] created Alibi, an open-source library for the generation of CFEs, and provided an open-source implementation of the loss function from Wachter et. al. [2017] (Klaise et al., 2021). They apply gradient descent to the objective loss function for correcting the distance between the initial counterfactual $\mathbf{x^{\prime}}$ and the original instance $\mathbf{x}$ (Klaise et al., 2021).
@@ 3.2 Segmented Sampling for Boundary Approximation Algorithm
Algorithm SSBA algorithm: Method of Boundary Point Generation
@@ 3.2.1 Algorithm Overview
In this section, we describe our central methodology, Segmented Sampling for Boundary Approximation (SSBA). Algorithm 1 is the core algorithm of SSBA, which generates a discrete set of decision boundary points for a binary classification problem. The algorithm takes in a trained model (model-agnostic), a dataset, labels, a threshold (number of total boundary points), and a precision value for the $\alpha$ used in the algorithm.
@@ 3.3 Ensuring feasibility compatible CF explanation method
Mahajan et. al. [2019] provide a method whereby they define additional example loss terms to their optimization. For unary constraints (single term constraints), they add a hinge loss $min(0,x_{v}^{{}^{\prime}}-x_{v})$ where $v$ selects a specific feature and superscript refers to the counterfactual of the instance $x$. For binary constraints, they use a linear model to measure the causal relationship between two features (Mahajan et al., 2019).
Mothilal, Sharma, and Tan [2020] allowed for two types of constraints in their implementation of DiCE. One constraint type allowed setting “feasible ranges for each [continuous] feature.” Another constraint type allows users to choose which variables or features can be changed while leaving others unchanged (Mothilal et al., 2020).
We let users define constraints in two different ways. In the first case, we allow users to define equality, less than, and greater than constraints for real-world constraints that constrain the choice of CFEs to fit reality. Mahajan et al. [2019] provide examples of unary constraints, such as $x_{age}^{cf}\geq x_{age}$, which restrict the choice of CFEs to either have the same age or older when compared to the age of the original instance (Mahajan et al., 2019).
@@ 4.1 Datasets
Secondly, the Adult Income dataset comprises over 26,000 fully labelled data points, each with an associated list of observational features and target values indicating whether the person’s annual income exceeds 50,000. We use DiCE’s variation of the dataset, which consists of $8$ features and one target value (Becker & Kohavi, 1996).
Thirdly, we use the heart disease dataset that consists of $13$ characteristics and one target value that indicates the presence or absence of heart disease. This dataset comprises over $500$ data points, with $13$ features representing various characteristics of the person, including age, sex, type of chest pain (cp), serum cholesterol levels, and maximum heart rate (thalach), among others (Janosi et al., 1989).
@@ 4.4 Visual Examples
5 Ablation Studies
Here, we study the effect of changing two key parameters with regard to the grid-based approach and the SSBA approach. These parameters are the resolution $R$ and the point epsilon $\epsilon$ for the grid-based method and the SSBA method, respectively.
The resolution size, $R$, is the size of the grid that we search through. For a given grid of size $R^{n}$ where $n$ is the number of features, we have to search through all $R^{n}$ points to find decision boundary points.
The point epsilon $\epsilon$ is the distance between two points from differing classes. In the case of the grid-based approach, this is the distance between two points that are predicted and labeled as part of separate classes. Similarly, in the case of the SSBA approach, this is the margin of error that we allow for two differing labeled and predicted points. This allows for the generated decision boundary point to lie within a small distance of two differing labeled points.
5.1 The Effect of Changing the Resolution
$R$ Boundary Points Found Runtime (s)
10 0 0.6
50 0 0.2
100 104 1.0
150 311 15.1
200 824 43.4
250 1612 108.7
Table 2: Boundary points and runtimes for varying $R$, the resolution size, for the grid-based approach where the grid is of size $R^{n}$ and where the number of features is $n=2$.
In Table $2$, we see the effect of changing the resolution size, $R$, for a logistic regression classifier applying grid-based approach. As we increase the overall resolution size, we increase the number of decision boundary points and the runtime. If we use a grid that is too sparse such as the case when $R=10$ or $R=50$, we don’t find any decision boundary points in the space between grid points. Given that we used a epsilon $\epsilon=0.1$, this means that we expect our grid points to be at most separated by a scalar distance of $0.1$. Otherwise, we don’t find any boundary points. However, as we increase the resolution up to $250$, we see that the number of boundary points start to increase, and we find $1612$ boundary points for a grid of size $250^{2}=62,500$ points.
5.2 The Effect of Changing the Point Epsilon
Method $\epsilon$ Prob. of class 1 Prob. of class 2
SSBA 10 $0.489$ $0.511$
SSBA 1 $0.491$ $0.509$
SSBA 0.1 $0.5001$ $0.49987$
SSBA 0.01 $0.50000$ $0.49999$
SSBA 0.001 $0.50000$ $0.49999$
Grid-based 10 $0.495$ $0.505$
Grid-based 1 $0.498$ $0.502$
Grid-based 0.1 $0.5003$ $0.4997$
Table 3: $\epsilon$ vs the average probabilities for both classes when applying a logistic regression classifier to generate decision boundary points.
In Table $3$, we find that decreasing $\epsilon$ in both the grid-based and SSBA approaches causes the probabilities for both classes to gradually approach $0.5$ or $50\%$ probability for each class of a logistic regression classifier and for a synthetic dataset of $2$ continuous features of equal class size. This makes sense for decision boundary points to have $50\%$ probabilities for both class since a logistic regression classifier uses a threshold of $50\%$ probability to differentiate points from two different classes.
This aligns with our expectation that as we reduce $\epsilon$, we get closer to the decision boundary since we are reducing the space between points of different classes, and thus we should expect that the probabilities should gradually approach $50\%$ for each class.
5.3 The Effect of Increasing $n$
Method $n$ $\epsilon$ Class 1 Class 2
SSBA 22 $10^{-4}$ $0.500$ $0.499$
SSBA $10$ $10^{-4}$ $0.500$ $0.499$
SSBA $50$ $10^{-4}$ $0.499$ $0.500$
SSBA $100$ $10^{-4}$ $0.5$ $0.5$
SSBA $1,000$ $10^{-4}$ $0.499$ $0.500$
SSBA $10,000$ $10^{-4}$ $0.500$ $0.499$
Table 4: $n$ vs the average probabilities for both classes when applying a logistic regression classifier to generate decision boundary points.
As we increase $n$, which is the size of the feature space, we observe that the model is very robust in high-dimensional spaces, capable of generating decision boundary points for a logistic regression classifier up to $\mathbf{R}^{10,000}$. The lack of a grid mainly explains this effect, as it requires an exponential number of additional points to find decision boundary points in the case of the grid-based approach and becomes harder to search.

### v4

@@ Abstract
What if a person never smoked? Would they have lung cancer? What identifies those who don’t have cancer from those who don’t? Counterfactuals are statements of “alternative hypothesis” that describe the outcomes of an alternative scenario (pearl2018book). They question the possibilities of a scenario that never occurred. While counterfactuals are hard to analyze, we can analyze counterfactual explanations which are more rigorously based. Counterfactual explanations define the “smallest” change to a data point such that the change flips the predicted outcome for a data point (molnar2025). In our article, we describe a method for generating counterfactual explanations in high-dimensional spaces using four steps that involve fitting our dataset to a model, finding the decision boundary, determining constraints on the problem, and computing the closest point (counterfactual explanation) from that boundary. We propose a discretized approach where we find many discrete points on the boundary and then identify the closest feasible counterfactual explanation. This method, which we later call Segmented Sampling for Boundary Approximation (SSBA), applies binary search to find decision boundary points, filters the decision boundary points for feasible points, and then returns the closest feasible point. Across four datasets of varying dimensionality, we show that our method can outperform current methods for counterfactual generation with reductions in distance between $5\%$ to $50\%$ in terms of the $L_{2}$ norm. Our method can also handle real-world constraints by restricting changes to immutable and categorical features, such as age, gender, sex, height, and other related characteristics such as the case for a health-based dataset. In terms of runtime, the SSBA algorithm generates decision boundary points on multiple orders of magnitude in the same given time when we compare to a grid-based approach. In general, our method provides a simple and effective model-agnostic method that can compute nearest feasible (i.e. realistic with constraints) counterfactual explanations. All of our results and our code can be found here at this link: https://github.com/ dsin85691/SSBA_For_Counterfactuals
@@ 1 Introduction
Imagine two scenarios: “If Joe had not smoked for $30$ years, he would have lived,” or “if Joe were 30 pounds lighter, he would not have heart disease.” While we may know that Joe smoked or was 30 pounds heavier, we may ask a question where we imagine outcomes to scenarios where the initial conditions were altered. These scenarios are among many described by Dr. Judea Pearl as “counterfactuals.” Counterfactuals are statements of “alternative hypothesis” that question the outcome given an alternative set of initial conditions (pearl2018book). When compared to other types of causal relationships commonly studied in machine learning, such as association or intervention, counterfactuals are “problematic” for study since we deal with outcomes that are not easily identifiable from the dataset. When given observational data for a person, we can never observe more than one potential outcome for a given person (pearl2018book). This makes it challenging to create logically consistent counterfactuals in a setting where we cannot perfectly replicate the same inputs. Standard machine learning methods, such as regression, can be helpful for questions of association where we observe how, for a set of variables $X$ and $Y$, a change in $X$ can alter the value of $Y$. With intervention, it is common to set up conditions so that other variables do not affect the primary variable of interest and its corresponding dependent variable. Randomized Control Trials (RCTs) were considered the “gold standard of causal analysis,” and it was common to randomize subjects into different groups to minimize the effect of confounding variables on treatment (pearl2018book). However, with counterfactuals, we must estimate the relationships between variables in a world where “observed facts are bluntly negated” (pearl2018book).
@@
To compute counterfactuals, Dr. Pearl relied on a method that made the use of structural causal models (SCMs) which are graphical models consisting of exogenous variables ($U$) and endogenous variables ($V$) and a “set of functions $f$ that assigns each variable in $V$ a value based on the other variables in the model” (pearl2016causal). Dr. Pearl provides a three-step process for computing any deterministic counterfactual with a structural causal model $M$, which includes abduction, action, and prediction (pearl2016causal). Abduction uses evidence $E=e$ to set the values of the exogenous variables in $M$. Action modifies the original model $M$ by removing the structural equations for the variables in $X$ and replacing them with the appropriate functions $X=x$ to obtain the model $M_{x}$ (pearl2016causal). Prediction uses the modified $M_{x}$ and the value of exogenous variables in $U$ to compute the value of $Y$, the consequence of the counterfactual (pearl2016causal). Similar to the idea used for Bayesian networks, we can use evidence or conditional information on the exogenous variables to simplify and construct a modified model that we can use to predict the outcome.
@@
In high-dimensional datasets, determining the structural causal model beforehand can be challenging. The use of causal discovery algorithms may require conditions such as causal sufficiency, which necessitates knowing all confounding variables within the dataset. Due to the difficulty of building a structural causal model and inferring causal relationships directly from the dataset, we address these issues by simplifying the problem to binary classification, assuming sufficient data for both classes, and relying on a formal definition of counterfactual explanations. As described by Dandl and Molnar [2025], a counterfactual explanation (CFE) describes “the smallest change to the feature values that changes the prediction to a predefined output” (molnar2025). Given the direct causal relationship between input and prediction of a machine learning model, we can observe “scenarios in which the prediction changes in a relevant way, like a flip in predicted class” (molnar2025). For CFEs, we find the smallest change across features such that the model flips a prediction value for a specific data point. For binary classification, such a flip would mean becoming a member of the other class as predicted by the model. In terms of criteria, we would like these explanations to be “as similar as possible to the instance regarding feature values” and require us to change as “few features as possible” (molnar2025). Finally, we would want feature values that are likely for the counterfactuals (molnar2025). Producing a CFE that generates illogical features, such as a younger person, or a person with an age greater than $150$, would not be realistic.
@@
To generate a CFE, we can employ multiple approaches, including randomly selecting counterfactuals, optimizing the choice of counterfactuals, and generating decision boundary points (our approach). One approach involves “randomly changing feature values of the instance of interest and stopping when the desired output is predicted” (molnar2025). As another approach, Wachter et. al. [2017] proposed an objective loss function where the two terms of the objective reduce the quadratic distance between the model prediction for the counterfactual $\mathbf{x^{\prime}}$ and a second term to minimize the distance between the instance $\mathbf{x}$ and the counterfactual $\mathbf{x^{\prime}}$ (molnar2025; wachter2017counterfactual). Across multiple iterations, we would find a counterfactual $x^{\prime}$ that minimizes the objective loss function, and then we return that counterfactual at the end.
@@
Dandl et. al. [2020] also implemented a four-objective loss function that minimized all of the above criteria for counterfactual explanations (molnar2025). Similar to Wachter et. al. [2017], they applied an combined objective loss function to optimize an appropriate counterfactual $\mathbf{x^{\prime}}$ selectively. $\mathbf{o_{1}},\cdots\mathbf{o_{4}}$ are objective functions that each individually represents one of the criteria above for CFEs.
@@
Among model-agnostic methods developed, Mothilal, Sharma, and Tan implemented DiCE (Diverse Counterfactual Explanations for ML), a Python package that generates CFEs for both gradient-based and model-agnostic methods (mothilal2020dice). Their gradient-based method optimizes over a combined loss function over all generated counterfactuals that take into account the difference between the machine predictions $f(c_{i})$ and the desired outcomes $y$ (first term), the distance between the counterfactual $c_{i}$ and the instance $x$, and the diversity among different counterfactuals.
@@
In their model-agnostic approaches, they adopt three different algorithms to generate a multitude of CFEs, which are their KDTrees algorithm, their random algorithm, and their genetic algorithm. The KDTree algorithm queries the $k$ closest counterfactuals from the dataset using KDTrees (mothilal2020dice). The random algorithm generates a random set of counterfactuals that are generally in the opposite class relative to the original instance (mothilal2020dice). Finally, the genetic algorithm tries to find the best counterfactuals close to the query point using a convergent algorithm that computes the best counterfactuals for a given generation based on a loss function taking into account criteria such as proximity and sparsity. It then generates a new generation based on $50\%$ of the fittest members of the current generation (mothilal2020dice). All of these methods use the original instance $\mathbf{x}$ and the counterfactual $\mathbf{x^{\prime}}$ to iterate and find a new counterfactual that matches the criteria and reduce the objective loss function between the counterfactual and the original instance.
@@ 2 Related Works
In this article, we primarily examine our methods with respect to DiCE and Alibi, which are both open-source Python packages for high-dimensional counterfactual generation. Wexler et al. [2019] present a similar methodology for nearest counterfactual generation. However, their approach selects a data point of a differing target class within the dataset (wexler2019if). While their tool can generate nearest counterfactual explanations based on the dataset, their “What-If” tool is specialized for two dimensions and does not readily apply to higher dimensions. Karimi et. al. [2019] implemented MACE (model-agnostic counterfactual explanation) with the use of SMT solvers, but their work applies a distance metric based on $L_{0}$, $L_{1}$, and $L_{\infty}$ and not $L_{2}$ for $n\geq 2$ (karimi2020model). Additionally, the discontinuation of the MACE project has made it difficult to replicate results applying $L_{2}$ norms. Due to limited research on counterfactual generation using $L_{2}$, we limit our work with respect to DiCE and Alibi.
@@ 3 Methodology
Our methodology is based on a discretized approach to find the nearest feasible counterfactual explanation.
@@
In the results section, we compare the computed counterfactual explanations and the runtime for various counterfactual generation methods which include DiCE, Alibi, and our SSBA method. As discussed earlier, Mothital et al. released DiCE, an open-source repository that adopts much of the recent research in the space of generating CFEs. Additionally, Janis Klaise et. al. [2021] created Alibi, an open-source library for the generation of CFEs, and provided an open-source implementation of the loss function from Wachter et. al. [2017] (JMLR:v22:21-0017). They apply gradient descent to the objective loss function for correcting the distance between the initial counterfactual $\mathbf{x^{\prime}}$ and the original instance $\mathbf{x}$ (JMLR:v22:21-0017).
@@ 3.2 Segmented Sampling for Boundary Approximation Algorithm
Algorithm Bisection Method Algorithm: Method of Boundary Point Generation
@@ 3.2.1 Algorithm Overview
In this section, we first apply the binary search method (bisection method) as suggested by Vlassopoulous et. al. (DBLP:journals/corr/abs-2006-07985). This generates a discrete set of decision boundary points for a binary classification problem. Our variant of the algorithm takes in a trained model (model-agnostic), a dataset, labels, a threshold (number of total boundary points), and a precision value for the $\alpha$ used in the algorithm.
@@ 3.3 Ensuring feasibility compatible CF explanation method
Mahajan et. al. [2019] provide a method whereby they define additional example loss terms to their optimization. For unary constraints (single term constraints), they add a hinge loss $min(0,x_{v}^{{}^{\prime}}-x_{v})$ where $v$ selects a specific feature and superscript refers to the counterfactual of the instance $x$. For binary constraints, they use a linear model to measure the causal relationship between two features (mahajan2019preserving).
Mothilal, Sharma, and Tan [2020] allowed for two types of constraints in their implementation of DiCE. One constraint type allowed setting “feasible ranges for each [continuous] feature.” Another constraint type allows users to choose which variables or features can be changed while leaving others unchanged (mothilal2020dice).
We let users define constraints in two different ways. In the first case, we allow users to define equality, less than, and greater than constraints for real-world constraints that constrain the choice of CFEs to fit reality. Mahajan et al. [2019] provide examples of unary constraints, such as $x_{age}^{cf}\geq x_{age}$, which restrict the choice of CFEs to either have the same age or older when compared to the age of the original instance (mahajan2019preserving).
@@ 4.1 Datasets
Secondly, the Adult Income dataset comprises over 26,000 fully labelled data points, each with an associated list of observational features and target values indicating whether the person’s annual income exceeds 50,000. We use DiCE’s variation of the dataset, which consists of $8$ features and one target value (adult_2).
Thirdly, we use the heart disease dataset that consists of $13$ characteristics and one target value that indicates the presence or absence of heart disease. This dataset comprises over $500$ data points, with $13$ features representing various characteristics of the person, including age, sex, type of chest pain (cp), serum cholesterol levels, and maximum heart rate (thalach), among others (heart_disease_45).